OpenStudy (anonymous):

an equation and two ordered pairs are given. Show that each pair is a solution of the equation. Then graph the two pairs to determine another solution. Y= 1/2x-1; (6,2), (0, −1)

OpenStudy (anonymous):

\[slope=\frac{-1-2}{0-6}=\frac{1}{2}\]

OpenStudy (anonymous):

-1=1/2(0)-1..........-1=-1 2=1/2(6)-1...........2=2

OpenStudy (anonymous):